) Harte Systems, Inc., a maker of electronic surveillance equipment, is considering selling to a well-known hardware chain the rights to market its home security system. The proposed deal calls for the hardware chain to pay Harte $30,000 and $25,000 at the end of years 1 and 2 and to make annual year-end payments of $15,000 in years 3 through 9. A final payment to Harte of $10,000 would be due at the end of year 10.

 i. Lay out the cash flows involved in the offer on a time line.

ii. If Harte applies a required rate of return of 12% to them, what is the present value of this series of payments?

 iii. A second company has offered Harte an immediate one-time payment of $100,000 for the rights to market the home security system. Which offer should Harte accept?                                              


1
Expert's answer
2021-03-04T10:36:57-0500

i.



ii.

"PV=\frac{FV}{(1+i)^1}+\frac{FV}{(1+i)^2}+...+\frac{FV}{(1+i)^n}=\frac{30000}{(1+0.12)^1}+\frac{25000}{(1+0.12^2}+...+\frac{10 000}{(1+i)^{10}}=104 508.28"




iii.

the current value from the first company is greater than the offer of the second. therefore, the offer of the second company is inappropriate